Shift report 20030219 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Feb 19 2003 - 23:46:37 EST

  • Next message: Apache: "Shift report 20030220 00:00-08:00"
    ------------------------------------------------------------
    Date: Wed Feb 19 23:46:36 2003, Shift: 16:00-24:00
    Supervisor: AlJ, with: 
    Log book#14 pages 26 - 27
    
    Shift Summary:
    ------------------------------------------------------------
    1. No beams during the shift. Maintenance up to 21.00
    2. TOFW-II mounted - 20 slats connected
    MRS put at 90 degrees. Dana Beavis asks data at this position to avoid possible interference between MRS and FS at forward angles.
    Magnets put at 1/4B (D1-D4), respectively, 1050B.
    3. CALFB, H1, C1 - some problems related to read from mainframe; they were fixed by H.Ito. No other problems.
    
    Run plan: 100K in FS at 12 degrees, trigger 6; 150 K in MRS at 90 degrees, trigger 3 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Wed Feb 19 2003 - 23:52:21 EST